OpenStudy (anonymous):

how do you solve this z^2=-12+16i for z in a +bi form ?

jimthompson5910 (jim_thompson5910):

-12+16i is in the form a+bi a = -12 b = 16 convert this to r*( cos(t) + i*sin(t) ) form where r = radius, t = angle theta r = sqrt( a^2 + b^2 ) r = sqrt( (-12)^2 + (16)^2 ) r = 20 t = arctan(b/a) t = arctan(16/(-12)) t = -53.1301

jimthompson5910 (jim_thompson5910):

so -12+16i converts to 20*( cos(-53.1301) + i*sin(-53.1301) ) and that's the same as 20*cis(-53.1301)

OpenStudy (anonymous):

thanks jim, you're the greatest B)

jimthompson5910 (jim_thompson5910):

using euler's formula, we can say 20*cis(-53.1301) = 20*e^(-53.1301i) so we now know this z^2=-12+16i is the same as z^2=20*e^(-53.1301i) now take the square root of both sides z = +-sqrt(20*e^(-53.1301i)) z = +-sqrt(20)*sqrt(e^(-53.1301i)) z = +-2*sqrt(5)*e^(-53.1301i/2) z = +-2*sqrt(5)*e^(-26.56505) then convert back to r*( cos(t) + i*sin(t) ) form

jimthompson5910 (jim_thompson5910):

z = +-2*sqrt(5)*e^(-26.56505) z = +-2*sqrt(5)*cis(-26.56505) z = +-2*sqrt(5)*[ cos(-26.56505) + i*sin(-26.56505) ] now convert to a+bi form z = +-2*sqrt(5)*[ cos(-26.56505) + i*sin(-26.56505) ] z = +-2*sqrt(5)*[ 0.894427 - 0.447214i] z = 2*sqrt(5)*[ 0.894427 - 0.447214i] or z = -2*sqrt(5)*[ 0.894427 - 0.447214i] z = 2*sqrt(5)*0.894427 - 2*sqrt(5)*0.447214i or z = -2*sqrt(5)* 0.894427 + 2*sqrt(5)*0.447214i z = 1.788854*sqrt(5) - 0.894428*sqrt(5)*i or z = -1.788854*sqrt(5) + 0.894428*sqrt(5)*i

jimthompson5910 (jim_thompson5910):

hold on, must have done something wrong, one sec

OpenStudy (anonymous):

i thin kyou just got it backwards! but thanks jim!

jimthompson5910 (jim_thompson5910):

no i found a better way, one sec

OpenStudy (anonymous):

alright :)

jimthompson5910 (jim_thompson5910):

Let z = a+bi z^2 = (a+bi)^2 z^2 = (a+bi)(a+bi) z^2 = a(a+bi)+bi(a+bi) z^2 = a^2 + abi + abi + b^2i^2 z^2 = a^2 + 2abi + b^2(-1) z^2 = a^2 + 2abi - b^2 z^2 = (a^2 - b^2) + (2ab)*i Real Part: a^2 - b^2 Imaginary Part: 2ab Real part in -12+16i is -12 so a^2 - b^2 = -12 Imaginary part is 16, so 2ab = 16 2ab = 16 ab = 8 b = 8/a ------------------ a^2 - b^2 = -12 a^2 - (8/a)^2 = -12 a^2 - 64/(a^2) = -12 q - 64/q = -12 ... let q = a^2 q^2 - 64 = 12q q^2 + 12q - 64 = 0 (q-4)(q+16) = 0 q-4 = 0 or q+16 = 0 q = 4 or q = -16 a^2 = 4 or a^2 = -16 a = 2, a = -2, a = 4i, a = -4i If a = 2, then b = 8/a b = 8/2 b = 4 So one solution is z = 2+4i If a = -2, then b = 8/a b = 8/(-2) b = -4 So another solution is z = -2-4i If you are restricting a and b to be real numbers, then those are the only two solutions

jimthompson5910 (jim_thompson5910):

So again, the only two solutions are 2+4i and -2-4i if you are restricting a and b to be real numbers

jimthompson5910 (jim_thompson5910):

yeah i did get the first part backwards, idk how, but i fixed it using this better approach